OpenStudy (anonymous):

by any chance do you know how to simp this? 2√ 3+√ 27-√ 1/3

OpenStudy (anonymous):

\[\sqrt{27}=\sqrt{9\times 3}=3\sqrt{3}\] might help

OpenStudy (anonymous):

collect like terms also \[\frac{1}{\sqrt{3}}=\frac{\sqrt{3}}{3}\]

OpenStudy (anonymous):

what would be the next step be? 5√3-√1/3. and am i even on the right track?

OpenStudy (swissgirl):

\(\large 5\sqrt{3}- {\sqrt{3} \over 3} \) \(\large {15\sqrt{3} \over 3} -{\sqrt{3} \over 3} \) \(\large {14\sqrt{3} \over 3}\)